OpenStudy (anonymous):

What is the value of X in 3/5 * X = 12?

OpenStudy (anonymous):

X = 5/3 X = 7 1/5 X = 20 X = 40

OpenStudy (anonymous):

2 Solve for X. X/4 = 2 1/4 *

OpenStudy (anonymous):

3 Solve for X. 5/6 * X = 30 *

OpenStudy (anonymous):

4 Solve for X. 1/2 * X = 9

OpenStudy (anonymous):

@DERRICKCURRY

OpenStudy (anonymous):

@Wendy.Ivette11714

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you help me with this

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i need to finish all of them like within 20 30 min :D so hopefully i understand

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the questions are up their ^ way up their

OpenStudy (anonymous):

/4(x-3)=2 doess this help 1/4x-3/4=2 1/4x=2 3/4 1/4x=11/4 x=11/4(4/1) x=11

OpenStudy (anonymous):

that is the first one yes?

OpenStudy (anonymous):

yuppers

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

x = 5/6 - 30 x = -29.167 alternatively, you may choose to solve it fractionally. x = 5/6 - 30 x = 5/6 - 180/6 x = -175/6

OpenStudy (anonymous):

wait i see

OpenStudy (anonymous):

x² + 9 (Not factorable under the reals, however it is factorable under the complex numbers) (x - 3i)(x + 3i), where i = √(-1) x² - 4 (x - 2)(x + 2) x² - 1 (x - 1)(x + 1)

OpenStudy (anonymous):

okay so... i understand that i just cant understand the answers

OpenStudy (anonymous):

it dose not mach to me

OpenStudy (anonymous):

ok hang on

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np

OpenStudy (anonymous):

whats the answer? so i can see what i am doing rong